OpenStudy (anonymous):

solve using the quadratic formula. 9x^2-15x+25=0

OpenStudy (anonymous):

I'm stuck on \[15 +- \sqrt{-15^{2}-900}/18\]

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\]\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[9x^2-15x+25=0\]\[x=\frac{15\pm\sqrt{(-15)^2-900}}{18}\]

OpenStudy (anonymous):

I have that down I just don't know how to simplify it.

OpenStudy (unklerhaukus):

\[(-15)^2=(-15)\times(-15)=\]

OpenStudy (anonymous):

225?

OpenStudy (unklerhaukus):

\[(-)(-)(15)(15)=(+)(5+10)(5+10)=(5\times5+2\times5\times10+10\times10)=255\]

OpenStudy (unklerhaukus):

yeah

OpenStudy (unklerhaukus):

so now what is 225-900?

OpenStudy (anonymous):

-675

OpenStudy (anonymous):

do i have to do\[\sqrt{-675} \]

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

[ sorry, i got disconnected]

OpenStudy (unklerhaukus):

That's right so far

OpenStudy (unklerhaukus):

When you have a parabola that is of this form\[y=ax^2+bx+c\] and and the discriminant \(\Delta=b^2−4ac\) There are three cases \(\Delta<0\) will not touch x axis - there will be no real solutions \(\Delta=0\) will touch x axis at one place only - one solution \(\Delta>0\) will touch x axis exactly twice - two solutions

OpenStudy (unklerhaukus):

the quadratic formula \[x=\frac{-b\pm\sqrt{\Delta}}{2a}\]

OpenStudy (unklerhaukus):

In your case the discriminant is negative, and the formula would have us take the square root of a negative number no real number is the square root of a negative. so there are no real solutions for x - [ this would give complex solutions but im not sure if you are looking for them ]

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

Are the complex solutions the same thing as imaginary solutions? I need imaginary solutions.

OpenStudy (anonymous):

Yes, they are the same thing.

OpenStudy (anonymous):

Do you know how I would get them for this problem?

OpenStudy (anonymous):

do you have \(\frac{ 15\pm \sqrt{-675} }{ 18 }\)? then you would want to simplify.

OpenStudy (anonymous):

Yes, that's exactly what I have right now. I'm not sure where to go from here.

OpenStudy (anonymous):

okay, I assume you're having trouble with simplifying \(\sqrt{-675}\) ? If so, first you want to break it up. I can rewrite it as \[\sqrt{-1}\sqrt{675} = \sqrt{-1}\sqrt{25}\sqrt{27}=\sqrt{-1}\sqrt{25}\sqrt{9}\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{25}=5 \] \[\sqrt{9}=3\] \[\sqrt{-1} = i \] right?

OpenStudy (anonymous):

Yes but don't forget you'll have the remaining \(\sqrt{3}\)

OpenStudy (anonymous):

How would I write that? Do I just multiply 5 and 3 and write \[\sqrt{3}i\] next to it?

OpenStudy (anonymous):

Yes, although I usually see i before the square root.

OpenStudy (anonymous):

So now I have\[x=15 \pm 15i \sqrt{3} / 18\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

That's not the answer, is it? Isn't there more to do?

OpenStudy (anonymous):

Is the answer different from the given answer? You can separate the +/- I guess

OpenStudy (anonymous):

No, it's not multiple choice. I just thought there was more to it. Thank you.

OpenStudy (unklerhaukus):

cancel the common factor of three

OpenStudy (unklerhaukus):

NB: the solutions are complex solutions. A number is complex when it has an imaginary And a real part,